OpenStudy (anonymous):

Graph the line 2x+3y=6.

OpenStudy (anonymous):

first put x = 0 and calculate the corresponding value of y say (x1, y1) next put y = 0 and calculate the corresponding value of x say (x2, y2) plot the two points and then draw a line passing through both the points

OpenStudy (anonymous):

So I have to make up my own x and y values?

OpenStudy (anonymous):

yes, as i suggested use zero for both once and get the corresponding value of other coordinate

OpenStudy (anonymous):

2x+3y=6. putting x = 0 2(0) + 3y = 6 3y = 6 y= 6/3 y= 2 so you get the point (0, 2)

OpenStudy (anonymous):

ok, you have the best response!

OpenStudy (anonymous):

ok what about 3x+5y=15

OpenStudy (anonymous):

follow same procedure.... first use x = 0 then y =0 then plot the two points and draw a line passing through them...